OpenStudy (anonymous):

Does 12 meters^3 = 15.696 yards^3?

OpenStudy (tkhunny):

Why not figure it out? \(15.696\;yd^{3}\cdot\left(\dfrac{3\;ft}{1\;yd}\right)^{3}\cdot\left(\dfrac{12\;in}{1\;ft}\right)^{3}\cdot\left(\dfrac{2.54\;cm}{1\;in}\right)^{3}\cdot\left(\dfrac{1\;m}{100\;cm}\right)^{3}\) Multiply by 1!

OpenStudy (anonymous):

Oh wow thank you! :)

OpenStudy (tkhunny):

?? What did you get?

OpenStudy (anonymous):

It was a multiple choice and 12 meters^3 does equal 15.696 yards^3. :)

OpenStudy (tkhunny):

Multiple choice? A) Yes, it's 12. B) No, it's not 12. Seriously?

OpenStudy (anonymous):

No, it had 3 others, I chose C. 3 centimeters^3 = 49.161 inches ^3

OpenStudy (anonymous):

and I got it right.